m 4 of 8 b la If Id g latd bte ctf -d -e -f g h ET f = -6, evaluate the determinant e h (A) 6 (B) -6 (C) -12 (D) 18
Answer #1

When a row of a determinant multiply by a constant,then determinant is also get multiplied by the same constant.

if one multiply of a row is added to another row ,then determinant will not change

given

\large \begin{vmatrix} a&b&c\\d&e&f\\g&h&i \end{vmatrix}=-6

add 2nd row to first row . This will not change determinant. So

\large \begin{vmatrix} a+d&b+e&c+f\\d&e&f\\g&h&i \end{vmatrix}=-6

multiply second row by -1 . Then determinant is also get multiplied by -1 . So

\large \begin{vmatrix} a+d&b+e&c+f\\-d&-e&-f\\g&h&i \end{vmatrix}=6

so answer is 6

option A